OpenStudy (anonymous):

I know I have to plug in the x but how without y? Question: Solve the system x + y = 6 and x - y = 2. What is the value of x? 6 2 4 0

OpenStudy (anonymous):

add

OpenStudy (anonymous):

\[x+y=6\\ x-y=2\] and in a column, the \(y\)'s go away and you get \[2x=8\] which tells you \(x=4\)

OpenStudy (anonymous):

not sure what you mean exactly by "I know I have to plug in the x" you cannot know what \(x\) is until you solve now you can find \(y\) by plugging in \(4\) for \(x\) in either equation

OpenStudy (anonymous):

i meant "add" in a column, not "and" sorry

OpenStudy (anonymous):

I see what you mean, im a visual learner so it takes time to read it and understand math, thanks!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

if you want a "visual" here is a picture of the two lines \(x+y=6\) and\(x-y=2\) you can see that they intersect at the point \((4,2)\) http://www.wolframalpha.com/input/?i=x%2By%3D6%2Cx-y%3D2

OpenStudy (anonymous):

thanks that link helped me!